Operations Report – April 15th

Crew 192 Operations Report 15Apr2018


Name of person filing report: Richard Blakeman

Non-nominal systems: None

Generator (hours run): 12 hours

Solar ‒ SOC 84%

Diesel – 75% fuel

Propane – 64% (telemetry reported by Shannon) psi.

Ethanol Free Gasoline (5 Gallon containers for ATV) – Approximately 4 gallons left in one and the other is empty

Water (trailer) – Empty, tank being cleaned in RAM.
Water (static) – 815 gallons (two external tanks)
Trailer to Static Pump used – No
Water (loft) – Static to Loft Pump used – Yes
Water Meter: 133514.4

Toilet tank emptied: Yes. Emptied today, 15 April.

ATV’s Used: Blue ATVs #1, #2

Oil Added? No but checked and OK.

ATV Fuel Used: 5%

# Hours the ATVs were Used today: 0.5

Notes on ATVs: All working well and ready for use, gas filled up today, all at 100%.

Deimos rover used: No
Hours: 119.8
Beginning charge: 100%
Ending charge: 100%
Currently charging: No

Sojourner rover used: ASSIGNED TO DIRECTOR
Beginning charge:
Ending charge:
Currently charging:

Spirit rover used: Yes
Hours: 38.2
Beginning charge: 100%
Ending charge: 98%
Currently charging: Yes

Opportunity rover used: No
Hours: 29.8
Beginning charge: 100%
Ending charge: 100%
Currently charging: Yes

Curiosity rover used: No
Hours: 25.9
Beginning charge: 100%
Ending charge: 100%
Currently charging: No

HabCar used and why, where? No

General notes and comments: A quiet day at the hab.

Summary of internet: 0MB at 1900.

Summary of suits and radios: Suits all charging, two one-piece suits and two two-piece suits used. Radios working fine. Spacesuit 9 ventilation wasn’t working yesterday, but worked today. May have just needed a charge/charging port may not have been inserted all the way. All good.

Summary of Hab operations: Stand-down day, crew relaxed around the hab. Also took the opportunity to perform some training. See commander’s report.

Summary of GreenHab operations: Plants watered twice.

Summary of ScienceDome operations: None

Summary of RAM operations: Trailer water tank prepped for additional cleaning in the RAM.

Summary of health and safety issues: None

Questions, concerns, and requests to Mission Support: None

Copyright © The Mars Society. All rights reserved. | Main Site